Ivanhoe Tire Co. just paid an annual dividend of $1.70 on its common shares. If Ivanhoe is expected to increase its annual dividend by 6.10 percent per year into the foreseeable future and the current price of Ivanhoe’s common shares is $16.40, what is the cost of common stock for Ivanhoe? (Round intermediate calculations to 4 decimal places, e.g. 0.1555 and final answer to 2 decimal places, e.g. 15.25%.)
Answer #1

Ans 17.10%

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
16.40 = 1.80370 / (Ke - 6.10%)
Ke - 6.10% = 1.80370 / 16.40
Ke - 6.10% = 11.00%
Ke = 17.10%
D1 = D0* (1 + g)
D1 = 1.70* (1 + 6.10%)
D1 = 1.80370